# Self-Alignment Radio Frequency Resonant Beam System for Information and Power Transfer

### v1

@@ I Introduction
To address these limitations and enhance high-efficiency transmission within the RF band, we propose the RF-based resonant beam system (RF-RBS). This system integrates RAAs and the resonance mechanism to facilitate self-alignment and achieve high-efficiency transmission for wireless power and information transfer. Fig. 1 illustrates a smart home scenario where the RF-RBS is applied, where the base station (BS) is mounted on the ceiling, allowing IoT and mobile devices to be wirelessly charged and communicate simultaneously. The RF-RBS enhances power transfer efficiency and communication reliability while improving radiation safety and reducing EM interference, contributing to the advancement of IoT technologies.
The contributions of this work are as follows
  • We propose an RF-RBS architecture using RAAs and the resonance mechanism, which can achieve watt-level power transmission with higher efficiency and supports high spectral efficiency downlink communication without the need for additional beam control mechanisms.
  • We developed an analytical model to explain the resonance mechanism based on EM propagation theory and self-reproduction model, enabling the numerical determination of the system’s status and performance.
  • We analytically characterize the system convergence behavior by taking into account the noise introduced by the RAAs and other components across iterations. Specifically, RF-RBS can converge within a limited number of iterations in the presence of noise. Once the effective operating range is exceeded, the introduced noise significantly reduces the system’s radiated power and transmission efficiency.
  • Compared to RD-BFS, RF-RBS achieves higher transmission efficiency due to its concentrated power distribution, resulting in up to a 16% improvement. Additionally, increasing the number of antennas in either the BS or the mobile target (MT) can extend the effective operating range of RF-RBS.
@@ II-A System Design of RF-RBS
The MT includes an antenna array of $M$ elements with phase-conjugate circuits. A power divider directs most energy to subsequent energy and communication parts, while the remaining signal is returned to the antenna through a phase-conjugate circuit. The output signal is rectified and delivered to the load in the energy processing unit. Information is extracted by demodulating the phase of the received signal.
@@ II-C Resonance Mechanism
II-D Multi-target Access
RF-RBS inherently supports broadcasting, allowing a single BS to form resonant beams with multiple MTs [41]. However, this broadcasting can lead to inefficient energy usage, as the system cannot simultaneously optimize power distribution for multiple targets. A Time Division Multiple Access (TDMA) scheme is employed within the RF-RBS to address this challenge.
In the TDMA scheme, time is divided into discrete slots, each allocated to a specific MT. During its designated time slot, the BS verifies the target’s activation status and establishes a stable connection. MTs not in their time slot remain inactive to prevent interference. The primary advantage of TDMA lies in its ability to manage power allocation and time distribution flexibly. MTs can be prioritized based on their criticality, with high-priority MTs receiving longer time slots and greater power, ensuring stable communication and energy transfer. Conversely, lower-priority MTs are allocated shorter slots with reduced power as necessary.
The analysis of the TDMA scheme for multi-target access in the optical RBC system is detailed in [41]. As with the optical RBC system, the MTs in RF-RBS can also suspend the resonant connection with the BS by turning off the reflection. By leveraging the TDMA scheme, RF-RBS can efficiently manage multi-target scenarios, ensuring fair power distribution and enhancing overall energy utilization. Therefore, the subsequent sections focus on the point-to-point analysis within RF-RBS.
@@ IV-A Parameter Setting
In the simulation, RF-BFS runs based on the model described above. Noise impacts the system’s convergence and stability and is also introduced as the initial perturbation signal to start the iteration process in RF-BFS. The RD-BFS simulation below uses a setup identical in size and antenna gain to the RF-RBS configuration. Instead of iteration, a central antenna of the MT is selected to emit an isotropic pilot signal. The BS is configured to record the power and phase of the received signals and then emit a phase-conjugate response. Since the radiated power of RD-BFS is configurable and lacks a comparable baseline, it is not displayed in the following simulation.
@@ IV-C Power Transmission and Efficiency Optimization
Figure 11 shows efficiency $\eta_{\text{cT}}$, BS radiated power $P_{\text{BS}}$ and DC output power $P_{\text{dc}}$ of RF-RBS and RD-BFS when MT moves along $z$-aixs with a distance of $\Delta z$. As $\Delta z$ increases, $\eta_{\text{cT}}$ for both systems initially remains stable but gradually decreases. The RF-RBS consistently outperforms the RD-BFS at medium to short distances, with a maximum efficiency improvement of 16% when the array size is $40\times 40$ and the distance is 1.8 m. Additionally, RF-RBS maintains a higher efficiency advantage over the effective distance, while RD-BFS’s efficiency declines faster. When the distance exceeds 3.4 m and 5.3 m, respectively, the efficiency of the RF-RBS drops to close to 0, where the transmission efficiency of the RF-RBS can no longer meet the gain-loss condition, and a stable resonance cannot be established. For radiated power $P_{\text{BS}}$, RF-RBS maintains the highest radiated power at short distances, which is determined by the PAs characteristic. As the distance increases, the radiated power decreases. After reaching the farthest working distance, the radiated power also drops to close to 0, where the radiated power is only the small power radiation under the influence of random noise. The DC power $P_{\text{dc}}$ shows a similar trend to $P_{\text{BS}}$ and maintains a higher conversion efficiency at shorter distances. The highest conversion efficiency exceeds 50%. And $P_{\text{dc}}$ is maintained at W-level within the working distance. Moreover, systems with more antenna elements achieve higher $\eta_{\text{cT}}$ values and can support longer transmission distances. Moreover, by comparing the efficiency changes in Fig. 11, with the increase of distance, the attenuation of RF-RBS is smoother, while that of RD-BFS is stepped. This is related to the fact that RD-BFS focuses only on the pilot antenna and ignores the other regions of MT, which causes the phase irregularity of the target region [25] and results in non-smooth attenuation of the efficiency.
@@ V Discussion
In the above, the point-to-point system is analyzed. In this section, we discuss the multi-target access within RF-RBS. The array composed of massive antennas can realize multi-direction transmission [xu2014MultiuserMISOBeamforming]. However, in RF-RBS, the direction and distance of MTs will affect the transmission loss, resulting in the multi-target scenario, each MT’s output power is uncontrollable [41]. To control the power separately, it is necessary to adopt an isolation scheme. In RF-RBS, the Time Division Multiple Access (TDMA) and Frequency Division Multiple Access (FDMA) schemes are feasible.
@@ V-A TDMA Scheme
In the TDMA scheme, time is divided into discrete slots, each allocated to a specific MT. During its designated time slot, the BS verifies the target’s activation status and establishes a stable connection. MTs not in their time slot remain inactive to prevent interference. The primary advantage of TDMA lies in its ability to manage power allocation and time distribution flexibly. MTs can be prioritized based on their criticality, with high-priority MTs receiving longer time slots and greater power, ensuring stable communication and energy transfer. Conversely, lower-priority MTs are allocated shorter slots with reduced power as necessary.
The analysis of the TDMA scheme for multi-target access in the optical RBC system is detailed in [41]. As with the optical RBC system, the MTs in RF-RBS can also suspend the resonant connection with the BS by turning off the reflection. By leveraging the TDMA scheme, RF-RBS can efficiently manage multi-target scenarios, ensuring fair power distribution and enhancing overall energy utilization. Therefore, the subsequent sections focus on the point-to-point analysis within RF-RBS.
The TDMA algorithm leverages a combination of Time Division Multiplexing (TDM) and Pulse Width Modulation (PWM), offering several advantages: 1. Multiple MTs can receive power simultaneously with varying charging intensities. 2. The BS can generate beams with consistent driving power for each receiver. 3. The driving power can be flexibly adjusted to meet specific requirements.
The devices within the work area, denoted as 1, 2, $\dots$, $N_{c}$, are managed using a digitized control scheme. The TDMA system divides the communication process into frames and slots, where each frame is divided into individual time slots allocated to each user. An additional time slot is reserved for new connection requests and the establishment of resonance.
For each user’s time slot, the time is divided into communication time ($T_{\text{com}}$) and resonance establishment time ($T_{\text{res}}$), which ensure the identification and activation of the target device and the creation of a stable resonant connection, respectively.
The total duration of each frame, $T_{\text{frame}}$, is calculated as:
$T_{\text{frame}}=\sum_{i=1}^{n_{t}}(T_{\text{com},i}+T_{\text{res},i})+T_{\text{req}},$
where $T_{\text{com},i}$ is the communication time, $T_{\text{res},i}$ is the resonance establishment time for the $i$-th user, and $T_{\text{req}}$ is the time for processing new connection requests.
During each user’s time slot, the BS first performs communication and device verification within $T_{\text{com},i}$, followed by $T_{\text{res},i}$ for establishing resonance. Data transmission only commences after these processes, with other MTs remaining idle to avoid interference.
Power allocation during each time slot follows the equation:
$P_{\text{total}}=\sum_{i=1}^{n_{t}}P_{i},$
where $P_{i}$ is the power allocated to the $i$-th user. The power allocation and slot duration can be adapted based on device priority. For example, critical devices can be assigned longer time slots and higher power to ensure stability, while less critical devices receive shorter slots and lower power.
V-A1 FDMA Scheme
In RF-RBS, the resonance establishment between the BS and MT need to align their operating frequencies, and the frequencies of LO sources in BS and MT define the their operating frequencies.
@@ V-A1 FDMA Scheme
To implement FDMA, the BS is divided into multiple sub-regions, each operating at a distinct frequency band. For example, in a 2x2 sub-array configuration, the BS can be divided into four sub-regions, each assigned with a different frequency. MTs tune their LO sources to match the frequency of the sub-region they intend to access. Multiple MTs establish resonant connections simultaneously and each works on a separate frequency.
The main advantage of FDMA is that each MT is assigned a dedicated frequency channel, minimizing interference. However, the maxinum number is limited. The total power available at the BS is distributed across the frequency bands based on device demand
$P_{\text{total}}=\sum_{j=1}^{n_{f}}P_{j}$ (26)
@@
where $n_{f}$ is the number of frequency bands, and $P_{j}$ is the power allocated to the $j$-th frequency band. By optimizing $P_{j}$ according to device requirements and channel conditions, FDMA efficiently supports multiple MTs.
@@ VI Conclusion
This paper demonstrated that RF-RBS operating in the RF band can achieve self-alignment and high-efficiency transmission based on the resonance mechanism. First, we introduced the RF-RBS architecture and the fundamentals of the RAA. Then, we established an analytical model to calculate power flow, resonance establishment, and communication performance. Numerical analysis revealed that the proposed RF-RBS consistently achieves self-alignment without beam control and provides higher transmission efficiency compared to RD-BFS, with improvements of up to 16%. Additionally, the results indicate the capability to transmit watt-level power and achieve 21 bps/Hz of downlink spectral efficiency in indoor settings. The more concentrated spatial power distribution of the RF-RBS enables high-efficiency transmission over long distances, highlighting the advantages of resonance in WPT and wireless communication applications. Future studies could investigate the system’s feasibility and performance under various scenarios and practical implementations. RF-RBS enables the IoT and mobile devices to operate without cables, presenting application prospects in smart homes, drones, and intelligent manufacturing. However, during implementation, several issues require further investigation. These include developing multiple safeguards for high-power transmissions, enhancing communication-perception integration, and integrating Reconfigurable Intelligent Surfaces (RIS) to expand the working range.

### v2

# Single-Frequency Self-Alignment RF Resonant Beam for Information and Power Transfer

@@ I Introduction
To address these limitations and enhance high-efficiency transmission within the RF band, we propose the RF-based resonant beam system (RF-RBS). This system integrates RAAs and the resonance mechanism to facilitate self-alignment and achieve high-efficiency transmission for wireless power and information transfer. Figure 1 illustrates a smart home scenario where the RF-RBS is applied, where the base station (BS) is mounted on the ceiling, allowing IoT and mobile devices to be wirelessly charged and communicate simultaneously. The RF-RBS enhances power transfer efficiency and communication reliability while improving radiation safety and reducing EM interference, contributing to the advancement of IoT technologies.
The contributions of this work are as follows:
  • We propose a single-frequency RF-RBS architecture based on the resonance mechanism produced by a pair of RAAs to achieve high-efficiency and watt-level power transmission and support high-spectral-efficiency communication, which do not need additional beam steering devices or algorithms.
  • We developed an analytical model to represent the resonance mechanism of RF-RBS based on EM propagation theory and self-reproduction model, enabling the numerical determination of the system’s status and performance.
  • We analytically characterize the system convergence behavior by taking into account the noise introduced by the RAAs and other components across iterations. Specifically, RF-RBS can converge within a limited number of iterations in the presence of noise. Once the effective operating range is exceeded, the system’s radiated power decrease significantly to ensure EM radiation safety.
  • Compared to RD-BFS, RF-RBS achieves higher transmission efficiency due to its concentrated power distribution, resulting in up to a 16% improvement. Additionally, increasing the number of antennas in either the BS or the mobile target (MT) can extend the effective operating range of RF-RBS without extending the convergence time required for the system to position and track the MT.
@@ II-A System Design of RF-RBS
The MT includes an antenna array of $M$ elements with phase-conjugate circuits. A power divider directs most energy to subsequent energy and communication parts, while the remaining signal is returned to the antenna through a phase-conjugate circuit. The output signal is then rectified, summed, and delivered to the load in the energy processing unit. Information is extracted by demodulating the phase of the received signal from the central element.
@@ IV-A Parameter Setting
In the simulation, RF-BFS operates based on the model described above. Noise impacts the system’s convergence and stability, which is also introduced as the initial perturbation signal to start the iteration process in RF-BFS. The RD-BFS simulation below uses a setup identical in size and antenna gain to the RF-RBS configuration. Instead of iterating, a central antenna of the MT is selected to emit an isotropic pilot signal. The BS is configured to record the power and phase of the received signals and then emit a phase-conjugate response [15]. Since the radiated power of RD-BFS is configurable and lacks a comparable baseline, it is not displayed in the simulation results.
@@ IV-C Power Transmission and Efficiency Optimization
Figure 11 shows efficiency $\eta_{\text{cT}}$, BS radiated power $P_{\text{BS}}$ and DC output power $P_{\text{dc}}$ of RF-RBS and RD-BFS when MT moves along $z$-aixs with a distance of $\Delta z$. As $\Delta z$ increases, $\eta_{\text{cT}}$ for both systems initially remains stable but gradually decreases. The RF-RBS consistently outperforms the RD-BFS at medium to short distances, with a maximum efficiency improvement of 16% when the array size is $40\times 40$ and the distance is 1.8 m. Additionally, RF-RBS maintains a higher efficiency advantage over the effective distance, while RD-BFS’s efficiency declines faster. When the distance exceeds 3.4 m and 5.3 m, respectively, the efficiency of the RF-RBS drops to close to 0, where the transmission efficiency of the RF-RBS can no longer meet the gain-loss condition, and a stable resonance cannot be established. For radiated power $P_{\text{BS}}$, RF-RBS maintains the highest radiated power at short distances, which is determined by the PAs characteristic. As the distance increases, the radiated power decreases. After reaching the farthest working distance, the radiated power also drops to close to 0, where the radiated power is only the small power radiation under the influence of random noise. The DC power $P_{\text{dc}}$ shows a similar trend to $P_{\text{BS}}$ and maintains a higher conversion efficiency at shorter distances. The highest conversion efficiency exceeds 50%. And $P_{\text{dc}}$ is maintained at W-level within the working distance. Moreover, systems with more antenna elements achieve higher $\eta_{\text{cT}}$ values and can support longer transmission distances. By comparing the efficiency changes in Fig. 11, with the increase of distance, the attenuation of RF-RBS is smoother, while that of RD-BFS is stepped. This is related to the fact that RD-BFS focuses only on the pilot antenna and ignores the other regions of MT, which causes the phase irregularity of the target region [25] and results in non-smooth attenuation of the efficiency.
@@ V Discussion
In the previous sections, we analyzed the point-to-point system. In this section, we discuss multi-target access within the RF-RBS. While an array composed of massive antennas can facilitate multi-directional transmission [54], the direction and distance of multiple MTs in RF-RBS can impact transmission loss, making power output to each MT uncontrollable [55]. To manage power outputs individually and enhance overall energy utilization, adopting an isolation scheme is essential. In RF-RBS, both Time Division Multiple Access (TDMA) and Frequency Division Multiple Access (FDMA) schemes are feasible.
@@ V-A TDMA Scheme
In RF-RBS, controlled resonance establishment is achievable by toggling the power return on or off in MT. Based on this mechanism, TDMA scheme can be used to achieve multi-target access. By leveraging TDMA, RF-RBS can efficiently manage scenarios involving multiple targets. In this scheme, time is segmented into discrete slots, each designated to a specific MT. During its allocated time slot, the BS verifies the target’s activation status and establishes a stable connection, while MTs outside their time slot remain inactive to avoid interference.
MTs within the BS’s operational area, denoted as 1, 2, $\dots$, $N_{c}$, are controlled digitally. Initially, the BS establishes connections with multiple MTs simultaneously, transmitting power and information through omnidirectional radiation, which leads to each MT sharing lower power. Subsequently, the BS allocates time slots to the MTs based on their priority and power needs.
The control in TDMA charging utilizes a digitalized rule comprising frames and slots. Each frame, denoted by $T_{f}$, is divided into individual time slots allocated to each MT. The allocation follows the rule that the sum of these MTs’ active durations must not exceed the frame length:
$\sum_{i=0}^{N_{c}}T_{\text{ON}_{i}}\leq T_{f},$ (26)
where $T_{\text{ON}_{i}}$ represents the active time of the $i$-th MT.
When $N_{c}$ MTs are present in the work area simultaneously, and they require connection to the BS for power transmission and communication, each MT is allocated time slots within a frame based on its needs and priority. Each MT’s time slot includes an additional resonance establishment period ($T_{\text{res}}$) to ensure the formation of a stable resonant connection. Therefore, the average power received by the $i$-th MT is calculated as:
$P_{\text{avg},i}=\frac{T_{\text{ON}_{i}}-T_{\text{res}}}{T_{f}}\eta_{c,i}P_{\text{BS}},$ (27)
where $\eta_{c,i}$ is the spatial transmission efficiency from the BS to the $i$-th MT, and $P_{\text{BS}}$ is the power radiated by the BS.
@@
The primary advantage of TDMA lies in its flexibility in managing power allocation and time distribution. MTs can be prioritized based on their criticality, with high-priority MTs receiving longer time slots and greater power to ensure stable communication and energy transfer. Conversely, lower-priority MTs are allocated shorter slots and reduced power as necessary. Thus, multiple MTs can receive power simultaneously with varying intensities. The BS can generate beams with consistent driving power for each MT, and the driving power can be flexibly adjusted to meet specific requirements.
V-B FDMA Scheme
In RF-RBS, the establishment of resonance between the BS and MTs requires the alignment of their operating frequencies, which are defined by the frequencies of the LO sources in both the BS and MTs. Utilizing this feature, multi-target access can be facilitated by assigning different operating frequencies to various MTs.
@@ V-B FDMA Scheme
In the FDMA scheme, the BS, equipped with massive antennas, is divided into multiple sub-regions, each operating at a distinct frequency band. Typically, in a $2\times 2$ sub-array configuration, the BS can be segmented into four sub-regions, each assigned with a different frequency. The MTs within the work area, denoted as 1, 2, $\dots$, $N_{c}$, tune their LO sources to match the frequency of the sub-region they intend to access. This arrangement allows multiple MTs to establish resonant connections simultaneously, with each operating on a separate frequency.
The total radiated power of the BS is distributed across these frequency bands based on the demand of the MTs:
$P_{\text{total}}=\sum_{i=1}^{N_{f}}P_{\text{BS}_{i}},$ (28)
@@
where $N_{f}$ represents the number of frequency bands, and $P_{\text{BS}_{i}}$ is the radiated power allocated to the $i$-th frequency band. By optimizing $P_{\text{BS}_{i}}$ according to the requirements of the MTs and channel conditions, FDMA can efficiently support multiple MTs.
The primary advantage of FDMA is that it assigns a dedicated frequency channel to each MT, minimizing interference. However, the maximum number of MTs that can be supported is limited by the number of available frequency bands in the BS. The integration of FDMA with the TDMA scheme can further enhance the capability of RF-RBS to support multi-target access.
@@ VI Conclusion
This paper demonstrated that RF-RBS operating in the RF band can achieve self-alignment and high-efficiency transmission based on the resonance mechanism. First, we introduced the RF-RBS architecture and the fundamentals of the RAA. Then, we established an analytical model to calculate power flow, resonance establishment, and communication performance. Numerical analysis revealed that the proposed RF-RBS consistently achieves self-alignment without beam control and provides higher transmission efficiency compared to RD-BFS, with improvements of up to 16%. Additionally, the results indicate the capability to transmit watt-level power and achieve 21 bps/Hz of downlink spectral efficiency in indoor settings. The more concentrated spatial power distribution of the RF-RBS enables high-efficiency transmission over long distances, highlighting the advantages of resonance in WPT and wireless communication applications. Future studies could investigate the system’s feasibility and performance under various scenarios and practical implementations. RF-RBS enables the IoT and mobile devices to operate without cables, presenting application prospects in smart homes, drones, and intelligent manufacturing. However, during implementation, several issues require further investigation, including i) developing multiple safeguards for high-power transmissions, ii) enhancing communication-perception integration, and iii) integrating Reconfigurable Intelligent Surfaces (RIS) to expand the working range.